# Slow dynamics and magnon bound states in the 2D long-range quantum Ising model

### v1

@@ Abstract
The dynamics of long-range quantum Ising models represents a current frontier in experimental physics, notably in trapped ions or Rydberg atomic systems. However, a theoretical description of these dynamics beyond 1D remains a significant challenge for conventional methods. Here, we address this challenge by means of neural quantum states to simulate global quenches from the fully polarized ferromagnetic state in the 2D quantum Ising model with power-law decaying interactions. From these numerically exact simulations, we find that the dynamics exhibit slow relaxation with long-lived oscillations. We explain this behavior through a theory for the formation of magnon bound states, which are generated, as we show, through effective attractive interactions between magnons that persist over several lattice sites due to the power-law nature of the interactions. Our results are readily observable in current quantum simulation platforms realizing long-range interacting models such as in Rydberg atomic systems.
Introduction.—Long-range interacting quantum systems arise across atomic, molecular, optical, and solid-state platforms [DefenuRMP2023], with prototypical realizations provided by spin models with power-law couplings in trapped ions [BrittonNature2012, RichermeNature2014, JurcevicNature2014, BohnetScience2016, ZhangNature2017, MonroeRMP2021, Kim2024], Rydberg atom arrays [BrowaeysNature2020, EbadiNature2021, BornetNature2023], and dipolar quantum gases [YanNature2013, HazzardPRL2014, ChristakisNature2023, PazPRL2013, AlaouiPRL2022]. Their tunable interaction range provides a controlled setting to investigate how nonlocal couplings reshape quantum dynamics.
In one dimension, theoretical and experimental studies have established that long-range interactions modify correlation spreading and the effective light-cone structure [Hauke2013PRL, RichermeNature2014], affect entanglement growth [JurcevicNature2014], and generate prethermal plateaus and long-lived oscillations that delay thermalization [LiuPRL2019]. These phenomena have been linked to emergent quasiparticles, such as domain walls or magnons, their mutual interactions, and the bound states they form [Bethe1931, Wortis1963, JurcevicNature2014, Kim2024, Kranzl2023, Fukuhara2013, Macri2021PRB]. However, extending these insights to two dimensions remains a central open problem. The exponential growth of the Hilbert space together with the rapid buildup of entanglement makes well-established methods such as exact diagonalization and standard tensor-network methods (e.g., tDMRG/MPS) challenging. Consequently, the quantum dynamics of 2D long-range models remain largely unexplored. The interplay between 2D geometry and power-law interactions raises new questions about the nature of the elementary excitations: Can nonlocal interactions in 2D bind distant excitations? How do they impact thermalization?
@@
In this work, we address these challenges by combining large-scale numerical simulations with an effective theoretical description. We employ neural quantum states (NQS) [CarleoScience2017, SchmittPRL2020, SchmittScience2022, schmitt2025] to simulate the real-time dynamics of the 2D long-range transverse-field Ising model following global quenches, accessing unprecedentedly large time scales and system sizes. Starting from the fully polarized ferromagnetic state, we observe a slow buildup of correlations accompanied by persistent oscillations. To interpret this behavior, we construct an effective few-body theory whose eigenspectrum shows excellent agreement with the Fourier spectral analysis of the time evolution. Crucially, the effective model reveals that the long-range interactions mediate a power-law attraction between magnons, yielding multi-magnon bound states even at separations of several lattice sites, leading to a large number of bound states. Thus, we identify these bound states as the relevant quasiparticles for the slow relaxation. These findings provide an experimentally accessible framework for exploring non-equilibrium dynamics in 2D long-range quantum matter.
@@
measured along the x-direction of the lattice. To compare dynamics across different $\alpha$ and $g$, we plot the normalized correlator $\tilde{C}(d,t)=C(d,t)\cdot\frac{8\mathcal{N}_{\alpha}J^{2}}{g^{2}}$, which removes the parameter-dependent amplitude scaling.
To reach system sizes beyond the limits of exact diagonalization for real-time simulations, we employ neural quantum states (NQS), a variational representation of quantum wavefunctions using artificial neural networks. Within NQS, the time-dependent state $|\psi_{\bm{\theta(t)}}\rangle=\sum_{S}\psi_{\bm{\theta(t)}}(S)|S\rangle$, with $|S\rangle$ denoting the full set of $2^{N}$ spin configurations, is represented through wave-function amplitudes $\psi_{\bm{\theta}(t)}(S)$ which are parametrized by network weights $\bm{\theta}$. The evolution of these parameters, $\bm{\theta}(t)$, is governed by the time-dependent variational principle. Importantly, NQS provides a numerically exact solution of the dynamics in the sense that convergence is guaranteed upon increasing the size of the underlying artificial neural network. Specifically, we use a convolutional neural network (CNN) architecture based on the ResNet framework Ref. [Kaiming2015, chen2024, chen2025], which is optimized for translational invariance and to capture 2D spatial correlations.
@@
Having established the structure of low-energy excitations in the pure Ising Hamiltonian, we now focus on the influence of the perturbation $V=-g\sum_{i}\hat{S}^{x}_{i}$. Since, for finite $\alpha$, the $\nu=0,1,2$ sectors together with the three-magnon sector containing at least a nearest-neighbor pair constitute well-separated excitation bands with gaps of $\mathcal{O}(J)$, the effect of a weak transverse field strength $g$ can be treated perturbatively. Applying a Schrieffer–Wolff (SWT) transformation up to second order in $g$ in each sector, we integrate out couplings to other sectors at leading order and thereby obtain an effective Hamiltonian acting solely within the sector with corrections to this structure arising only at order $\mathcal{O}(g^{3}/J^{2})$. For the two-magnon manifold, imposing translational symmetry and restricting to the zero-momentum subspace, where both the initial state and the dynamics reside, yields the effective Hamiltonian
@@
with $\delta U(\mathbf{d})$ denoting corrections proportional to $g^{2}/J$, which at large distances $|\mathbf{d}|$ provide only a subleading contribution to the attractive interactions proportional to $\frac{-g^{2}}{8J\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{2(\alpha-1)}}$, as long as $\alpha>2$. The off-diagonal amplitudes $t_{\mathbf{d},\mathbf{d}^{\prime}}$ are real and symmetric, describing correlated long-range pair hoppings, which at large $|\mathbf{d}-\mathbf{d^{\prime}}|$ are given by $t_{\mathbf{d},\mathbf{d}^{\prime}}\approx\frac{-g^{2}}{8J\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}-\mathbf{d}^{\prime}|^{\alpha}}$ (explicit derivations are provided in the supplementary material). For all numerical calculations, however, we employ the full (non-asymptotic) effective Hamiltonian derived via the Schrieffer–Wolff transformation. Notice that while an attractive interaction potential might be naturally associated with the formation of bound states in the case of short-range kinetics [yang1989, chadan2003], we find that the interplay between the long-range attractive potential and the correlated pair long-range hoppings produces a rich spectrum of two-magnon bound states.
@@
To achieve a more refined frequency resolution, the inset displays the FFT of the exact dynamics for a $5\times 5$ system evolved up to $Jt=200$. The resulting spectrum agrees very well with the effective predictions, except for the red line. This deviation arises because, for $L\leq 5$, the excitations above the two-magnon manifold, those involving three or more magnons do not form a well-separated band, which limits the applicability of the SWT description. Nevertheless, exact diagonalization reveals that this peak originates from a compact four-magnon cluster. These observations show that the slow dynamics are controlled by a sparse set of well-defined few-magnon excitations rather than a broad many-body continuum. Furthermore, the clear one-to-one correspondence between the spectral peaks extracted from the NQS dynamics and the energy gaps of the effective theory confirms the accuracy of both approaches up to at least $g/J=0.5$.
Emergence of magnon bound states.—Having validated the few-body theory with our exact numerical simulations with NQS, we now aim to analyze in more detail the two-magnon eigenstates. In this section, we address three central questions: (i) how the spatial structure of the bound states depends on the interaction exponent $\alpha$, (ii) whether the long-range attractive potential favors bound states beyond the nearest-neighbor separation, and (iii) how these bound states influence the dynamics. In the NN case ($\alpha=\infty$), only nearest-neighbor bound pairs exist, with all other two-magnon configurations forming a broad scattering continuum [Tindall2024PRL, Luka2025PRB]. Long-range interactions qualitatively change this picture: the attractive potential extends over many lattice spacings, potentially stabilizing bound states at separations far beyond nearest neighbors.
To characterize these states, we analyze the eigenstates $|\psi(\mathbf{d})\rangle$ of the effective two-magnon Hamiltonian $\hat{H}_{\text{eff}}^{(2)}$. For each eigenstate, we compute its inverse participation ratio (IPR), $\mathrm{IPR}[\psi]=\sum_{\mathbf{d}}|\psi(\mathbf{d})|^{4}$, which distinguishes localized bound pairs (high IPR) from delocalized scattering states (low IPR). Intermediate IPR values correspond to quasilocalized states, whose wavefunctions contain a localized core accompanied by a weak delocalized tail. We also compute the mean pair separation, $\bar{d}=\sum_{\mathbf{d}}|\psi(\mathbf{d})|^{2}\,|\mathbf{d}|$, to determine the average relative positions. Figure 3 shows the resulting IPR-$\bar{d}$ distribution for several power-law exponents $\alpha$ on a $101\times 101$ lattice with $g/J=0.2$, together with representative real-space probability densities in the relative coordinate $(\mathbf{r}_{1}-\mathbf{r}_{2})$.
@@
Discussion and Outlook. We study the dynamics of the 2D long-range transverse-field Ising model using large-scale neural quantum state simulations, accessing system sizes and evolution times beyond exact methods. We observe slow relaxation with persistent oscillations after a global quench from a ferromagnetically ordered state. We show that these dynamics are governed by a set of two-magnon bound states stabilized by the long-range attractive interaction. Future experiments with Rydberg arrays and trapped ions could directly probe these spatially extended bound magnons via site-resolved correlation spectroscopy, providing a real-time visualization of bound-pair formation and decay.
Beyond the specific transverse-field Ising dynamics considered here, our framework applies more broadly to long-range interacting systems with different types of kinetic terms—whether generated by bosonic hopping or exchange interactions. The essential ingredient responsible for the emergence of multi-magnon bound states is the long-range attractive potential. Our effective theory reveals that the spin-flip kinetic term renormalizes into a long-range hopping process. Notably, bound states remain stable despite this long-range kinetic motion. Thus we expect that such extended bound states are a generic and robust characteristic of 2D long-range quantum matter.
Data availability.—Data to reproduce all figures are available at Zenodo [dataset_long-range_quantum_Ising_model].
@@
Acknowledgements.—The NQS dynamics is simulated using the jVMC package [Schmitt2022], and the exact dynamics are performed with the QuSpin package [Weinberg2017]. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 853443). This work was supported by the German Research Foundation (DFG) via project 492547816 (TRR 360). We gratefully acknowledge the scientific support and high-performance computing resources provided by the LiCCA HPC cluster at the University of Augsburg and by the Erlangen National High Performance Computing Center (NHR) of the Friedrich-Alexander-Universität Erlangen–Nürnberg (NHR Project No. nqsQuMat). The LiCCA cluster is co-funded by the DFG (Project ID 499211671). NHR funding is provided by federal and Bavarian state authorities, and the NHR@FAU hardware is partially funded by the DFG (Grant No. 440719683).
@@
Note added.—During the preparation of this manuscript, we became aware of related work by Kaltenmark et al. [kaltenmark2025], who study spectroscopic signatures in a kinetically constrained long-range interacting two-dimensional spin system. While their microscopic model and parameter regime differ from ours, both works highlight the emergence of nontrivial elementary excitations in long-range interacting quantum systems.
End Matter
Appendix A: Neural quantum state simulations.—We represent the many-body wavefunction by a neural quantum state (NQS) $\psi_{\bm{\theta}}(S,t)$, where $S\!\in\!\{\pm 1\}^{L\times L}$ and $\bm{\theta}$ are trainable parameters. Real-time dynamics follows the time-dependent variational principle (TDVP), which projects the Schrödinger equation onto the NQS manifold:
$\sum_{b}S_{ab}\,\dot{\theta}_{b}\;$ $=\;-i\,F_{a},$ (A1)
$\qquad S_{ab}=\langle O_{a}^{\dagger}O_{b}\rangle_{c},$ $\quad F_{a}=\langle O_{a}^{\dagger}E_{\rm loc}\rangle_{c},$ (A2)
with logarithmic derivatives $O_{a}(S,t)=\partial_{\theta_{a}}\ln\psi_{\bm{\theta}}(S,t)$, connected correlators $\langle\cdot\rangle_{c}$ over $|\psi_{\bm{\theta}}|^{2}$, and local energy $E_{\rm loc}(S)=\sum_{S^{\prime}}H_{SS^{\prime}}\,\psi_{\bm{\theta}}(S^{\prime})/\psi_{\bm{\theta}}(S)$. We integrate the equation of motion $S\,\dot{\bm{\theta}}=-i\mathbf{F}$ using a Heun solver, which provides a second-order update for $\bm{\theta}(t)$.
We employ a convolutional neural network (CNN) based on the ResNet framework, with Gaussian Error Linear Unit (GELU) activation functions. We enforce lattice translation invariance by using the convolutions with periodic padding. The convolutional blocks produce a final set of real-valued feature maps. These maps are then split to generate the real and imaginary parts of an intermediate complex feature map. A global logsumexp pooling operation is applied across all spatial positions and feature channels of this map to produce a base scalar log-wavefunction. To produce the final NQS, we additionally impose the point-group symmetries of the square lattice (rotations and reflections), by averaging the exponentiated base wavefunction over all symmetric operations at the output layer. We use networks consisting of two ResNet blocks with $3\times 3$ filter kernels and 16 feature channels for all simulations presented in this work.
All simulations are performed in the $\hat{S}^{z}$ basis, where each configuration $S$ in $\psi_{\bm{\theta}}(S)$ denotes a spin configuration along the $z$-direction. Although the Hamiltonian 1 is written in the $ZZ+X$ form in the main text, we implement the NQS simulations using the equivalent rotated $XX+Z$ representation. This choice ensures that the ferromagnetically ordered initial state corresponds to a uniformly distributed wavefunction in the configuration space, rather than a single delta peak, which removes instabilities coming from zeroes of the wavefunction.
Expectation values in Eq. (A2) are estimated using Markov-chain Monte Carlo sampling with Metropolis single-spin-flip updates. For each time step, we use $5\times L^{2}$ sweeps for thermal equilibration, followed by $3\times 10^{4}-4\times 10^{4}$ measurement sweeps. We use an adaptive Heun scheme for the integration of the TDVP equation, with integration tolerance of $10^{-4}$.
Appendix B: Spectral analysis.— To extract excitation energies, we compute the discrete Fourier transform (FFT) of $\mathcal{O}(t)\in\{C(d,t),\,\langle\hat{S}_{i}^{z}(t)\rangle\}$ over a finite time and apply a Hamming window to reduce spectral leakage. With $t_{n}=n\,\Delta t$ and $n=0,\ldots,N_{t}\!-\!1$, we define $w_{n}=0.54-0.46\cos\!\left(\tfrac{2\pi n}{N_{t}-1}\right)$, and the transform
$\mathcal{F}({\mathcal{O}}(t))(\omega_{k})=\sum_{n=0}^{N_{t}-1}w_{n}\,\big[\mathcal{O}(t_{n})-\overline{\mathcal{O}}\big]\,e^{i\omega_{k}t_{n}},$ (A3)
where $\overline{\mathcal{O}}=\tfrac{1}{N_{t}}\sum_{n=0}^{N_{t}-1}\mathcal{O}(t_{n})$ removes the zero-frequency component and $\omega_{k}=\tfrac{2\pi k}{N_{t}\,\Delta t}$ with $k=0,\ldots,N_{t}\!-\!1$. Peaks in $|\mathcal{F}({\mathcal{O}}(t))(\omega_{k})|$ identify the frequencies governing the dynamics. These peak positions are compared to energy differences $|E_{i}^{\nu}-E_{j}^{\nu^{\prime}}|$ obtained by diagonalizing the effective few-magnon Hamiltonians $\hat{H}_{\mathrm{eff}}^{(\nu)}$, allowing us to assign each spectral line to a single-magnon or a multi-magnon bound state.
@@
Supplemental Material: Slow dynamics and magnon bound states in the 2D long-range quantum Ising model
Appendix A Schrieffer–Wolff transformation.
@@
$\hat{H}_{\rm eff}^{(\nu)}=\sum_{m}E_{\nu}(m)\,|m\rangle\!\langle m|+\frac{g^{2}}{2}\sum_{m,n,\beta}\langle m|\hat{T}|\beta\rangle\,\langle\beta|\hat{T}|n\rangle\left(\frac{1}{E_{m}-E_{\beta}}+\frac{1}{E_{n}-E_{\beta}}\right)|m\rangle\!\langle n|,$ (S1)

### v2

# Persistent coherent quantum dynamics in 2D long-range magnets via magnon binding

@@ Abstract
The dynamics of 2D long-range quantum magnets represents a current frontier in experimental physics such as in Rydberg atomic systems or trapped ions. In this work we address theoretical challenges in understanding these dynamics by combining large-scale neural quantum state simulations with an effective theory. Our findings uncover a mechanism for persistent coherent quantum dynamics and slow relaxation in 2D long-range quantum magnets. Demonstrated on the 2D transverse-field quantum Ising model with power-law decaying interactions, we observe long-lived oscillatory behavior after quenching the system from a ferromagnetic product state. We explain this phenomenon by the formation of magnon bound states, generated by effective attractive long-range magnon interactions. Our results highlight a generic mechanism for long-lived quantum coherence in 2D quantum magnets that can be directly observed in current quantum simulation platforms.
Introduction.
@@
Long-range quantum magnets arise across atomic, molecular, optical, and solid-state platforms [15], with prototypical realizations provided by spin models with power-law couplings in trapped ions [7, 30, 21, 5, 41, 27, 23], Rydberg atom arrays [8, 16, 6], and dipolar quantum gases [38, 19, 13, 14, 1]. In 1D, theoretical and experimental studies have established that long-range interactions modify correlation spreading and the effective light-cone structure [18, 30], affect entanglement growth [21], and generate prethermal plateaus and long-lived oscillations that delay thermalization [25]. These phenomena have been linked to emergent quasiparticles, such as domain walls or magnons, their mutual interactions, and the bound states they form [4, 37, 21, 23, 24, 17, 26].
However, extending these insights to two dimensions remains a central open problem. The exponential growth of the Hilbert space together with the rapid buildup of entanglement makes well-established methods such as exact diagonalization and standard tensor-network methods (e.g., tDMRG/MPS) challenging. Consequently, the quantum dynamics of 2D long-range magnets remain largely unexplored. Beyond the numerical challenge, the interplay of 2D geometry and power-law interactions can give rise to dynamical mechanisms that are distinct from both the one-dimensional case and the nearest-neighbor limit, raising basic questions: What are the relevant elementary excitations? Can nonlocal interactions in 2D bind excitations over several lattice sites? How does this impact thermalization?
In this Letter, we address these questions by combining large-scale neural quantum state (NQS) simulations [9, 31, 33, 32] with an effective theoretical description. We uncover a generic mechanism for persistent coherent quantum dynamics and slow relaxation in 2D long-range quantum magnets. As an experimentally relevant benchmark, we focus on the 2D transverse-field Ising model with power-law interactions, accessing system sizes and evolution times beyond established numerical approaches. Quenching from a fully polarized ferromagnetic product state, we observe slow relaxation characterized by persistent, long-lived oscillations, a striking departure from the general expectation of rapid thermalization. To interpret this behavior, we construct an effective theory whose eigenspectrum shows excellent agreement with the Fourier spectral analysis of the real-time dynamics. Crucially, the effective model reveals that long-range interactions mediate a power-law attraction between magnons, yielding multi-magnon bound states even at separations of several lattice sites and producing a large manifold of bound states. We thus identify extended magnon bound states as the quasiparticles governing slow relaxation in the 2D long-range Ising model, and argue that magnons form bound states whenever long-range attractive interactions compete with weak kinetic terms, making magnon binding a generic mechanism for persistent coherent dynamics in 2D long-range magnets.
@@
measured along the x-direction of the lattice. For comparison across different $\alpha$ and $g$, we plot the rescaled correlator $\tilde{C}(d,t)=C(d,t)\cdot\frac{8\mathcal{N}_{\alpha}J^{2}}{g^{2}}$, which provides a suitable normalization across multiple parameter regimes.
To reach system sizes beyond the limits of exact diagonalization for real-time simulations, we employ neural quantum states (NQS), a variational representation of quantum wavefunctions using artificial neural networks. Within NQS, the time-dependent state $|\psi_{\bm{\theta(t)}}\rangle=\sum_{S}\psi_{\bm{\theta(t)}}(S)|S\rangle$, with $|S\rangle$ denoting the full set of $2^{N}$ spin configurations, is represented through wave-function amplitudes $\psi_{\bm{\theta}(t)}(S)$ which are parametrized by network weights $\bm{\theta}$. The evolution of these parameters, $\bm{\theta}(t)$, is governed by the time-dependent variational principle. Importantly, NQS provides a numerically exact solution of the dynamics in the sense that convergence is guaranteed upon increasing the size of the underlying artificial neural network. Specifically, we use a convolutional neural network (CNN) architecture based on the ResNet framework Ref. [20, 11, 12], which is optimized for translational invariance and to capture 2D spatial correlations.
@@
Having established the structure of low-energy excitations in the pure Ising Hamiltonian, we now focus on the influence of the perturbation $V=-g\sum_{i}\hat{S}^{x}_{i}$. Since, for finite $\alpha$, the $\nu=0,1,2$ sectors together with the three-magnon sector containing at least a nearest-neighbor pair constitute well-separated excitation bands with gaps of $\mathcal{O}(J)$, the effect of a weak transverse field strength $g$ can be treated perturbatively. Applying a Schrieffer–Wolff (SWT) transformation up to second order in $g$ in each sector, we integrate out couplings to other sectors at leading order and thereby obtain an effective Hamiltonian acting solely within the sector with corrections to this structure arising only at order $\mathcal{O}(g^{3}/J^{2})$. For the two-magnon manifold, imposing translational symmetry and restricting to the zero-momentum subspace, yields the effective Hamiltonian
@@
with $\delta U(\mathbf{d})$ denoting corrections proportional to $g^{2}/J$, which at large distances $|\mathbf{d}|$ provide only a subleading contribution to the attractive interactions proportional to $\frac{-g^{2}}{8J\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}|^{2(\alpha-1)}}$, as long as $\alpha>2$. The off-diagonal amplitudes $t_{\mathbf{d},\mathbf{d}^{\prime}}$ are real and symmetric, describing correlated long-range pair hoppings, which at large $|\mathbf{d}-\mathbf{d^{\prime}}|$ are given by $t_{\mathbf{d},\mathbf{d}^{\prime}}\approx\frac{-g^{2}}{8J\mathcal{N}_{\alpha}}\frac{1}{|\mathbf{d}-\mathbf{d}^{\prime}|^{\alpha}}$ (explicit derivations are provided in the supplementary material). For all numerical calculations, however, we employ the full (non-asymptotic) effective Hamiltonian derived via the Schrieffer–Wolff transformation. Notice that while an attractive interaction potential might be naturally associated with the formation of bound states in the case of short-range kinetics [39, 10], we find that the interplay between the long-range attractive potential and the correlated pair long-range hoppings produces a rich spectrum of two-magnon bound states.
@@
To achieve a more refined frequency resolution, the inset displays the FFT of the exact dynamics for a $5\times 5$ system evolved up to $Jt=200$. The resulting spectrum agrees very well with the effective predictions, except for the red line. This deviation arises because, for $L\leq 5$, the excitations above the two-magnon manifold, those involving three or more magnons, do not form a well-separated band, which limits the applicability of the SWT description. Nevertheless, exact diagonalization reveals that this peak originates from a compact four-magnon cluster. These observations show that the slow dynamics are controlled by a sparse set of well-defined few-magnon excitations rather than a broad many-body continuum. Furthermore, the clear one-to-one correspondence between the spectral peaks extracted from the NQS dynamics and the energy gaps of the effective theory confirms the accuracy of both approaches up to at least $g/J=0.5$.
We emphasize that the effective Hamiltonian is broadly applicable to 2D long-range magnets. Whenever a microscopic model features power-law decaying ferromagnetic couplings, the low-energy sector above a fully polarized ferromagnetic state is governed by few-magnon excitations. Adding kinetic terms such as spin flips, hopping, or pair creation and annihilation, or any combination thereof, yields, under a SW transformation, an effective Hamiltonian of very similar structure.
Emergence of magnon bound states.—Having validated the effective theory with our exact numerical simulations with NQS, we now aim to analyze in more detail the two-magnon eigenstates. In this section, we address two central questions: (i) whether the long-range attractive potential binds two magnons beyond the nearest-neighbor separation, (ii) how the spatial structure of these states depends on the interaction exponent $\alpha$. To characterize these states, we analyze the eigenstates $|\psi(\mathbf{d})\rangle$ of the effective two-magnon Hamiltonian $\hat{H}_{\text{eff}}^{(2)}$. For each eigenstate, we compute its inverse participation ratio (IPR), $\mathrm{IPR}[\psi]=\sum_{\mathbf{d}}|\psi(\mathbf{d})|^{4}$, which distinguishes localized bound pairs (high IPR) from delocalized scattering states (low IPR). Intermediate IPR values correspond to quasilocalized states, whose wavefunctions contain a localized core accompanied by a weak delocalized tail. We also compute the mean pair separation, $\bar{d}=\sum_{\mathbf{d}}|\psi(\mathbf{d})|^{2}\,|\mathbf{d}|$, to determine the average relative positions. Figure 3 shows the resulting IPR-$\bar{d}$ distribution for several power-law exponents $\alpha$ on a $101\times 101$ lattice with $g/J=0.2$, together with representative real-space probability densities in the relative coordinate $(\mathbf{r}_{1}-\mathbf{r}_{2})$.
@@
Discussion and Outlook. Our NQS simulations and effective theory establish magnon binding as a generic microscopic mechanism for persistent coherent dynamics and slow relaxation in 2D long-range magnets. Future experiments with Rydberg arrays and trapped ions could directly probe these spatially extended bound magnons using site-resolved correlation spectroscopy, enabling real-time visualization of bound-pair formation and decay. This mechanism is fundamentally distinct from the 1D case, where domain walls are the elementary excitations and long-range interactions confine domain-wall kinks [25], and from the 2D nearest-neighbor limit, where slow dynamics has been linked to approximate Hilbert-space fragmentation due to the near-conservation of domain-wall length in the weak-field regime [40, 2, 3, 35, 29]. Beyond the transverse-field Ising model studied here, our framework, in principle, applies broadly to 2D long-range magnets with diverse kinetic processes. Accordingly, we expect that magnon binding could be a generic feature of 2D long-range quantum magnets, potentially also in 3D, where magnons again constitute the elementary excitations.
Data availability.—Data to reproduce all figures are available at Zenodo [28].
@@
Acknowledgements.—The NQS dynamics is simulated using the jVMC package [34], and the exact dynamics are performed with the QuSpin package [36]. This project has received funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 853443). This work was supported by the German Research Foundation (DFG) via project 492547816 (TRR 360). We gratefully acknowledge the scientific support and high-performance computing resources provided by the LiCCA HPC cluster at the University of Augsburg and by the Erlangen National High Performance Computing Center (NHR) of the Friedrich-Alexander-Universität Erlangen–Nürnberg (NHR Project No. nqsQuMat). The LiCCA cluster is co-funded by the DFG (Project ID 499211671). NHR funding is provided by federal and Bavarian state authorities, and the NHR@FAU hardware is partially funded by the DFG (Grant No. 440719683).
@@
Note added.—During the preparation of this manuscript, we became aware of related work by Kaltenmark et al. [22], who study spectroscopic signatures in a kinetically constrained long-range interacting two-dimensional spin system. While their microscopic model and parameter regime differ from ours, both works highlight the emergence of nontrivial elementary excitations in long-range interacting quantum systems.
@@
Supplemental Material: Persistent coherent quantum dynamics in 2D long-range magnets via magnon binding
Appendix A Schrieffer-Wolff transformation.
@@
$\hat{H}_{\rm eff}^{(\nu)}=\sum_{m}E_{\nu}(m)\,|m\rangle\!\langle m|+\frac{g^{2}}{2}\sum_{m,n,\beta}\langle m|\hat{T}|\beta\rangle\,\langle\beta|\hat{T}|n\rangle\left(\frac{1}{E_{m}-E_{\beta}}+\frac{1}{E_{n}-E_{\beta}}\right)|m\rangle\!\langle n|,$